OpenStudy (anonymous):

How would I find the center and radius for: x^2+(y+4)^2=9

hero (hero):

The general equation of a circle has the form \((x - h)^2 + (y - k)^2 = r^2\) where \((h,k)\) is the center of the circle and \(r\) is the radius.

OpenStudy (anonymous):

what if there is no h?

hero (hero):

(x - 0)^2 = x^2 h is an implied zero

OpenStudy (anonymous):

ok and does 9 have to be a square root?